OpenStudy (anonymous):

given the functions F(x)=1-3x^2 and g(x)=square root of 4-x. find fog

OpenStudy (anonymous):



OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f(\sqrt{4-x})=1-3\sqrt{4-x}^2\] then algebra

OpenStudy (anonymous):

okay can you do out the algebra i did it wrong? and what is domain of fog